OpenStudy (anonymous):

log(x+4) = log x + log 4

OpenStudy (ryan123345):

Hey there! I see that you are new to OpenStudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! If you have any additional questions involving the website feel free to ask the people with the purples A's next to their name. They are the ambassadors and will answer all your questions regarding the site. And also read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Have a great day and I hope you have fun!

OpenStudy (mertsj):

log a+log b=log ab Use that on the right side.

OpenStudy (anonymous):

log (x+4) = log (x times 4)

OpenStudy (anonymous):

that is where I then go blank

OpenStudy (raden):

x times x = 4x then cancel out the "log" from both sides

OpenStudy (mertsj):

If the logs are equal and the bases are the same, then the arguments must be the same. x+4=4x

OpenStudy (anonymous):

Okay. I wasn't sure that I could cancel the logs. So, x = 4/3

OpenStudy (raden):

though one :P

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

Thank you!

OpenStudy (mertsj):

yw